What ratio of CH3NH2 to CH3NH3* is needed to prepare a pH 10.00 buffer? (Kb for CH3NH2 is 4.4 % 10-4)
Answer #1

mixture of amine and its salt act as basic buffer

pOH = pKb + log [CH3NH3+] / [CH3NH2]

pKb = - log Kb = - log [4.4 x 10-4] = 3.36

pH = 10.00

pOH = 14 - 10.00

pOH = 4.00

4.00 = 3.36 +  log [CH3NH3+] / [CH3NH2]  

log [CH3NH3+] / [CH3NH2] = 0.64

[CH3NH3+] / [CH3NH2] = 4.36

[CH3NH2] / [CH3NH3+] = 0.23